1.given the following information
Pieces of pizza total Utility marginal utility
Qx TUx MUx
0 0
2 8
4 14
6 18
8 20
10 20
12 18

Required:
A. Compute for MUx. Show your solution
B. Based on the table , how many pieces of pizza the consumer will ,consume to maximize his satisfaction
2. Given the following information
Price of burger(Pb)=P50
Price of spaghetti (Ps)= P75
MU of burger(MUb)=180 utils
MU of spaghetti(MUs)= 225 utils
Required:
A. Compute for the consumers equilibrium
B. What is the decision of the consumers ? Will he buy burger or spaghetti ? Justify your answr.
1
Expert's answer
2021-07-21T15:02:33-0400

1)

A)The marginal utility can be calculated as the ratio of change in total utility to the change in quantity.

"marginal \space utility=\frac{Change\space in \space total\space utility}{Change\space in\space Quantity}"

When the quantity increased from 0 to 2, the total utility increased from 0 to 8. Therefore the marginal utility of the first quantity is calculated as,

"Marginal \space utility=\frac{8−0}{2−0}\\=\frac{8}{2}\\=4"

Similarly, the marginal utility of the other quantities is calculated and given in the table below.



B)

The table shows that the marginal utility is positive till the 8th unit is consumed and it is zero for the tenth unit. After 10 units are consumed, the marginal utility is negative. That, the amount utility gained by the consumer is increasing till the 8th unit, although at a decreasing rate. After the 10th unit, the utility is decreasing. Therefore the consumer gets the maximum satisfaction at the 10th unit.


2)A. The consumer will be in equilibrium when the ratio of marginal utility of the two goods and services is equal to ratio of price of the both goods and services. 

"\frac{MU_b}{MU_s} = \frac{P_b}{P_s}"

where MU

b

MUb shows the marginal utility of burgers, MU

s

MUs shows the marginal utility of spaghettis 

Putting the value of these, we'll get

"\frac{180}{225 } >\frac{50}{75}\\0.8> 0.66"

This shows that the consumer is not in equilibrium as the ratio are not equal. 


B. The consumer would decide to increase the consumption of spaghettis and decrease the consumption of burgers. This will decrease the marginal utility of burgers and increase the marginal utility of spaghettis and help to restore the equilibrium. 

In conclusion, the consumer would want to be in equilibrium where the equality of the ratio persist. In order to be there, he would buy more spaghettis than burgers.